# thanks: Corresponding author

### v2

@@ II.7 Summary of the SSM
[…]
(b) Particle: PS, Target: aluminum
(c) Particle: PMMA, Target: aluminum
Figure 6: (a) Schematic of the particle impact charging apparatus (not to scale). For each reference impact condition, at least one hundred fresh, uncharged particles impact the target under identical conditions. (b) Reference impact results for PS particles. Histograms display the raw measured charge distributions for two particle sizes. According to the SSM, the distribution of the larger particles is predicted by scaling the statistical parameters obtained from the smaller particles. (c) Reference impact results for PMMA particles.
@@ III.1 Statistical parameters: $\mu_{0}$, $\sigma_{0}$, $\gamma_{0}$, $\Delta Q_{0,\mathrm{min}}$
Table 1: Reference impact experiments.
Particle PS PS PMMA
Target aluminum aluminum aluminum
[…]
$v_{n}$ (m/s) $1.1$ $1.1$ $0.9$
Impacts $241$ $317$ $100$
$\mu_{0}$ (fC) 68.5-68.5 306.6-306.6 4.1-4.1
$\sigma_{0}$ (fC) $72.3$ $183.5$ $6.5$
$\gamma_{0}$ (-) 0.56-0.56 0.45-0.45 0.19-0.19
$\Delta Q_{0,\mathrm{min}}$ (fC) 462.0-462.0 1842.0-1842.0 21.9-21.9
@@
The reference impact conditions and measured statistical parameters are listed in Table 1 for two powders, polystyrene (PS) and polymethyl methacrylate (PMMA), interacting with an aluminum wall. Experimental charge distributions and model fits are shown in Figs. 6b and 6c.
Figure 7 plots the sensitivity of the statistical parameters with the number of impacts for $r=538$ µm sized PS particles. According to the figure, the statistical errors drop below 20% after 150 impacts.
@@ III.2 Test of the scaling laws
For each material pair, the model requires only a single reference condition. Nonetheless, we conducted two reference experiments for PS to evaluate the accuracy of the model’s scaling. These cases differ solely in particle size. As shown in Fig. 6b, the SSM successfully predicts the impact statistics of the larger particles based on the smaller-particle reference, with errors between 13% and 36%.
@@ IV CFD simulations
Figures 9b and 9c present the corresponding charge distributions of the 200 µm PS and PMMA particles at selected time instances. For both materials, the SSM captures the smooth transition of the distributions from centering around uncharged particles toward peaking near the saturation charge.
@@ IV.3 Size-dependent bipolar charging
Interestingly, the saturation charge of the particles depends not only on their size but, in contrast to the simulations shown in Fig. 9a, also on the particle material. Here, the saturation of particle charge is not governed by the charge transfer from the wall to the particle and the associated decline of $\alpha_{w}/\alpha_{0}$. Instead, in the periodical box simulations (Fig. 12), charging terminates when all charging sites on all particles are depleted. Because the particles follow random collision histories and thus accumulate charge at different rates, the final charge is not uniform across particles (as in Figs. 9b and 9c) but follows a distribution. The final charge distribution for PS particles at $t=47.4$ s is shown in Fig. 12c.
@@ V.1 Physical assumptions of the SSM
A central simplification (Fig. 2) is the separation of the impact charge between two surfaces into two independent charge transfers in opposite directions: A deterministic transfer from a conductor and a stochastic transfer from an insulator. The experiments of Matsuyama et al. [12] support this assumption. They showed that the impact charge of a conductive particle on a conductive target varies very little, whereas the impact charge from an insulating particle exhibits strong variability.
@@
As a stochastic model, the SSM does not aim to predict the charge evolution of individual particles but rather the evolution of the charge distributions of the entire powder. Accordingly, it does not attempt to resolve mechanisms that depend on the history of individual particles [14] but captures powder-wide charge distributions resulting from such mechanisms’ aggregation. We use a different particle in each of the $M\geq 100$ reference impact experiments per powder to capture these global effects.
@@ Appendix C Variable impact charge PMMA
For completeness, Fig. 14 presents the variable impact charge of PMMA particles, analogous to Fig. 10 for PS. Figures 10 and 14 show qualitatively similar behavior.

### v3

@@ II.2 Scaling the statistical parameters: $\mu_{w}$, $\mu_{i}$, $\sigma_{i}$, $\gamma_{i}$
Using Eqs. 11d, 11c and 11b for particle-particle charge transfer implies that the charge-transfer distribution for same-material contacts can be inferred from the distribution measured for different-material contacts. We test this assumption in Section III.2.
@@ III Reference impact experiment
[…]
Figure 6: (a) Schematic of the particle impact charging apparatus (not to scale). For each reference impact condition, fresh, uncharged particles impact the target under identical conditions.
(a) Particle: PS, Target: aluminum
(b) Particle: PMMA, Target: aluminum
[…]
Figure 7: Results of the reference impact experiment. The histograms display the raw measured charge distributions; the dashed curves the predictions by SSM. (a) PS particles of two different sizes. The SSM predicts the distribution of the larger particles by scaling the statistical parameters obtained from the smaller particles. (b) Reference impact results for PMMA particles. (c) Identical PMMA particles impacting an aluminum and a PMMA target. The SSM predicts the distribution of the same-material impacts by scaling the statistical parameters obtained from the different material impacts.
Table 1: Reference impact experiments.
Particle PS PS PMMA PMMA PMMA
Target aluminum aluminum aluminum aluminum PMMA
[…]
$v_{0}$ (m/s) 1.1 1.1 0.9 1.7 1.7
Impacts 241 317 100 46 165
$\mu_{0}$ (fC) -68.5 -306.6 -4.1 -444.9 56.7
$\sigma_{0}$ (fC) 72.3 183.5 6.5 121.3 191.9
$\gamma_{0}$ (-) -0.56 -0.45 -0.19 -0.03 0.02
$\Delta Q_{0,\mathrm{min}}$ (fC) -462.0 -1842.0 -21.9 -825.6 -778.8
@@ III.1 Statistical parameters: $\mu_{0}$, $\sigma_{0}$, $\gamma_{0}$, $\Delta Q_{0,\mathrm{min}}$
The reference impact conditions and measured statistical parameters are listed in Table 1 for two powders, polystyrene (PS) and polymethyl methacrylate (PMMA), interacting with an aluminum wall. Experimental charge distributions and model fits are shown in Fig. 7.
Figure 8 plots the sensitivity of the statistical parameters with the number of impacts for $r=538$ µm sized PS particles. According to the figure, the statistical errors drop below 20% after 150 impacts. For each reference experiment in Fig. 7, we measured impacts until the statistical parameters converged.
@@ III.2 Test of the scaling laws
For each material pair, the model requires only a single reference condition. Nonetheless, we conducted two reference experiments for PS to evaluate the accuracy of the model’s scaling. These cases differ solely in particle size. As shown in Fig. 7a, the SSM successfully predicts the impact statistics of the larger particles based on the smaller-particle reference. The absolute error of the scaling by SSM compared to the measured distribution is 0.040 pC for the mean, 0.041 pC for the standard deviation, and 0.73 for the skewness.
@@
Further, we test the assumption that the charge-transfer distribution for same-material contacts can be obtained by scaling the distribution measured for different-material contacts. To this end, we conducted two reference experiments with PMMA particles, differing only in the target material, which was aluminum in one case and PMMA in the other. As shown in Fig. 7c, for the investigated PMMA particles, the SSM successfully predicts the statistical characteristics of same-material impacts based solely on the statistics of different-material charging. The absolute errors between the SSM-scaled predictions and the measured distributions are 0.057 pC for the mean, 0.020 pC for the standard deviation, and 0.02 for the skewness.
@@ IV CFD simulations
Figures 10b and 10c present the corresponding charge distributions of the 200 µm PS and PMMA particles at selected time instances. For both materials, the SSM captures the smooth transition of the distributions from centering around uncharged particles toward peaking near the saturation charge.
@@ IV.3 Size-dependent bipolar charging
Interestingly, the saturation charge of the particles depends not only on their size but, in contrast to the simulations shown in Fig. 10a, also on the particle material. Here, the saturation of particle charge is not governed by the charge transfer from the wall to the particle and the associated decline of $\alpha_{w}/\alpha_{0}$. Instead, in the periodical box simulations (Fig. 13), charging terminates when all charging sites on all particles are depleted. Because the particles follow random collision histories and thus accumulate charge at different rates, the final charge is not uniform across particles (as in Figs. 10b and 10c) but follows a distribution. The final charge distribution for PS particles at $t=47.4$ s is shown in Fig. 13c.
@@ V.1 Physical assumptions of the SSM
A central simplification (Fig. 2) is the separation of the impact charge between two surfaces into two independent charge transfers in opposite directions: A deterministic transfer from a conductor and a stochastic transfer from an insulator. The experiments of Matsuyama et al. [12] support this assumption. They showed that the impact charge of a conductive particle on a conductive target varies very little, whereas the impact charge from an insulating particle exhibits strong variability. The assumption that the charge-transfer distribution for same-material impacts can be derived from the stochastic component of the different-material distributions is, for the PMMA particles under consideration, validated in Fig. 7c.
@@
As a stochastic model, the SSM does not aim to predict the charge evolution of individual particles but rather the evolution of the charge distributions of the entire powder. Accordingly, it does not attempt to resolve mechanisms that depend on the history of individual particles [14] but captures powder-wide charge distributions resulting from such mechanisms’ aggregation. We use a different particle in each of the total 869 reference impact experiments to capture these global effects.
@@ Appendix C Variable impact charge PMMA
For completeness, Fig. 15 presents the variable impact charge of PMMA particles, analogous to Fig. 11 for PS. Figures 11 and 15 show qualitatively similar behavior.